OpenStudy (anonymous):

The scale on a map is approx. 5/8 inches to 15 miles. If the distance from city A to city B is 9.5 inches on the map, how far apart are the two cities?

OpenStudy (anonymous):

\[\frac{5}{8}=0.625\] \[\frac{.625}{15}=\frac{9.5}{x}\]

OpenStudy (anonymous):